# Labeling eigenstates of qubit-cavity systems based on the continuity of qubit occupancy: Detecting resonances to higher excited qubit states

### v2

@@ IV How to choose appropriate energy window
By recalling the results presented in Sec. III, a resonant peak of the cavity frequency is replaced with a sudden drop in failed labeling. Based on this empirical rule, the labeling with the energy window $\delta=1.5\times 10^{-2}\hbar\omega_{c}$ should be appropriate. This expectation can be verified by comparing the trajectory obtained by the dynamics under cavity driving. The red-dotted line in Fig. 4(b) is the trajectory in the $\braket{\hat{N}_{q}}$-$\braket{\hat{c}^{\dagger}\hat{c}}$ plane under the cavity driving with $E/\omega_{c}=1.0\times 10^{-2}$ and $\omega_{d}/\omega_{c}=1.001$. We use the Fock state of the cavity up to |$\ket{500}_{c}$ for the numerical simulation of the dynamics. The trajectory obtained by the dynamics follows the photon number dependence of the qubit occupancy with the energy window $\delta/\hbar\omega_{c}=1.5\times 10^{-2}$. The dynamics ensure that the labeling with $\delta/\hbar\omega_{c}=1.5\times 10^{-2}$ is appropriate. This rough matching with the dynamics also implies that a drop of the cavity frequency around $n\sim 240$ will not be replaced with a resonant peak.
@@
Through this example, one can find the importance of using some energy windows and comparing those results. From the obtained results, one can eliminate labeling with an evident discontinuity in physical quantities. When a sudden change in the cavity frequency is replaced with a peak in another labeling, the labeling with the sudden change would be wrong. This empirical rule is useful for choosing an appropriate energy window without accessing dynamics. For the numerically expensive but definite confirmation, one can compare the results of labeling to those obtained from dynamics.
V Offset charge dependence of resonance
@@
As we have already observed, the indication of the resonance to higher excited states can be seen in the photon-number dependence of the cavity frequency obtained by the labeling methods. Such resonance can lead to leakage from the computational space [19, 23] and to chaotic dynamics of the transmon [24, 21, 28]. The identification of its location would be an essential in designing the superconducting qubits. Since the labeling method proposed in this study can be applied to a broader parameter region compared to the previous method [21], the proposed method is appropriate for the identification.
@@
In this section, we concentrate our attention on the offset charge $N_{g}$ dependence of the resonant points. The offset charge contains contributions from charge impurities. One of the important characteristics of the transmon superconducting qubit is robustness of its resonant frequency to the offset charge noise. In other words, the energy difference between the first excited and the ground states has little dependence on the offset charge in the transmon regime $E_{J}/E_{C}\gg 1$. It should be noted that the eigenenergies of higher excited states depend on the offset charge [12]. Consequently, the locations of resonances to the higher excited states are expected to depend on the offset charge noise.
@@
Figure 5: Photon-number dependence of the cavity frequency with the offset charges from 00 to $0.5$ for (a) the ground-state ladder |$\ket{\overline{g},\overline{n}}$ and (b) the excited-state ladder |$\ket{\overline{e},\overline{n}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2})$ for the system parameters. For the cases $N_{g}=0.1$ in the ground-state ladder and $N_{g}=0.3$ in the excited-state ladder, the energy window $\delta$ is set to $1.5\times 10^{-2}\hbar\omega_{c}$. For the other cases, we set $\delta$ to $1.0\times 10^{-2}\hbar\omega_{c}$.
@@ VI Summary
In this paper, we proposed a new method to label the eigenstates of a qubit-cavity system based on the continuity of qubit occupancy. Starting from the eigenstate without photon occupation, the system is expected to follow the labeled states under adiabatic injection of photons into the cavity. Besides, the effective resonant frequency of the cavity with photon occupation can be estimated from the eigenenergies of the labeled states. Our proposed method finds candidates for the next labeled state based on their eigenenergies. Then, one selects the next labeled state that makes the change of the qubit occupancy as small as possible. Comparing the obtained photon-number dependence of the cavity frequency and the evolution under cavity driving, we confirmed that our proposed scheme can be applied to a broader parameter range compared to the previous approach [21].
With the proposed method, we investigated the offset charge dependence of the cavity frequency in the transmon-cavity system. From the photon-number dependence of the cavity frequency, one can estimate the presence of the resonances to the higher transmon excited states which can lead to the leakage effects from the computational space [19, 23] and the chaotic dynamics of the transmon [21, 24, 28]. In the system we studied, the location of the resonance fluctuates depending on the offset charge. Even though the qubit resonant frequency of the transmon is robust to the charge noises, the resonance to higher excited states is not. Without controlling the offset charge, it is possible to encounter a resonance to higher excited states with a relatively small photon number $n\sim 10$.
The photon-number dependence of the cavity frequency and the location of the resonance to higher excited states are essential information for designing superconducting qubits. Our proposed method is applicable to other superconducting qubits such as the fluxonium [13] or qubits with $d$–wave superconductors [15, 16]. Their robustness to the charge noise, including effects from higher excited states, can be evaluated with a labeling method. A stable labeling method is essential in assessing the robustness of existing or new high-performance superconducting qubits.

### v3

@@ III Comparison of labeling methods
By recalling the results presented in Sect. III, a resonant peak of the cavity frequency is replaced with a sudden drop in a labeling following the resonant transition. Based on this empirical rule, the labeling with the energy window $\delta=1.5\times 10^{-2}\hbar\omega_{c}$ should be appropriate to evaluate the effective cavity frequency with avoiding resonant peaks. This expectation can be verified by comparing the trajectory obtained by the dynamics under cavity driving. The red-dotted line in Fig. 4(b) is the trajectory in the $\braket{\hat{N}_{q}}$-$\braket{\hat{c}^{\dagger}\hat{c}}$ plane under the cavity driving with $E/\omega_{c}=1.0\times 10^{-2}$ and $\omega_{d}/\omega_{c}=1.001$. We use the Fock state of the cavity up to |$\ket{500}_{c}$ for the numerical simulation of the dynamics. The trajectory obtained by the dynamics follows the photon number dependence of the qubit occupancy with the energy window $\delta/\hbar\omega_{c}=1.5\times 10^{-2}$. The dynamics ensure that the labeling with $\delta/\hbar\omega_{c}=1.5\times 10^{-2}$ is appropriate to avoid resonant transitions. This rough matching with the dynamics also implies that a drop of the cavity frequency around $n\sim 240$ will not be replaced with a resonant peak.
When two labelings with energy windows $\delta_{1}$ and $\delta_{2}$ show different behaviors at a resonant point, the level repulsion between resonant states can be bounded by using the parameters $\delta_{1}$ and $\delta_{2}$. The two labelings with energy windows $\delta_{1}$ and $\delta_{2}$ are supposed to give the same label up to the photon number $n-1$. Then, the next eigenenergies $\varepsilon_{p,n}$ are assumed to be different in the two labelings, and we represent them as $E_{r}+$(--)$\Delta/2$ for $\delta_{1}$($\delta_{2}$). Here, $E_{r}$ is the resonant energy and $\Delta$ is the level repulsion emerged from the interaction between resonant states. We choose the energy window for the larger eigenenergy as $\delta_{1}$. Consequently, the level repulsion is positive, $\Delta>0$. In such situations, the energy window conditions $|\varepsilon_{p,n}-\varepsilon^{\prime}_{p,n}|\leq\delta/2$ for the two labelings can be rewritten as
$2(\varepsilon^{\prime}_{p,n}-E_{r})-\delta_{1}$ $\leq\Delta\leq 2(\varepsilon^{\prime}_{p,n}-E_{r})+\delta_{1}$ (12)
$-2(\varepsilon^{\prime}_{p,n}-E_{r})-\delta_{2}$ $\leq\Delta\leq-2(\varepsilon^{\prime}_{p,n}-E_{r})+\delta_{2}.$ (13)
We note that a candidate for eigenenergy $\varepsilon^{\prime}_{p,n}$ is the same for the two labelings up to $n$. Since $\Delta$ is positive, the inequality $-2(\varepsilon^{\prime}_{p,n}-E_{r})<\delta_{1}$ is satisfied from Eq. (12). With the inequality and Eq. (13), one can bound the level repulsion $\Delta$ as
$0<\Delta<\delta_{1}+\delta_{2}.$ (14)
Especially when the parameters $\delta_{1}$ and $\delta_{2}$ have the same order of magnitude $\delta_{1}\sim\delta_{2}\sim\delta$, the bound for the repulsion is given as
$0<\Delta\lesssim 2\delta.$ (15)
Thus, the energy window parameter $\delta$ roughly bounds the allowable size of level repulsions that the proposed labeling method can follow. To change labeling behaviors at a resonant point, the parameter $\delta$ should be varied by the same magnitude as $\Delta$.
@@
Through the example presented in this section, one can find the importance of using some energy windows and comparing those results. When a sudden change in the cavity frequency is replaced with a peak in another labeling, the labeling with the sudden change would follow a resonant transition. This empirical rule is useful for choosing an appropriate energy window to avoid resonant transitions without accessing dynamics. For the numerically expensive but definite confirmation, one can compare the results of labeling to those obtained from dynamics.
V Offset-charge dependence of resonance
@@
As we have already observed, the indication of the resonance to higher excited states can be seen in the photon-number dependence of the cavity frequency obtained by the labeling methods. Such resonance can lead to leakage from the computational space Sank et al. (2016); Khezri et al. (2023); Shillito et al. (2022). The identification of its location would be an essential in designing the superconducting qubits. Since the labeling method proposed in this study can be applied to a broader situation compared to the previous method Shillito et al. (2022), the proposed method is appropriate for the identification.
@@
In this section, we concentrate our attention on the offset-charge $N_{g}$ dependence of the resonant points. The offset charge contains contributions from charge impurities. One of the important characteristics of the transmon superconducting qubit is robustness of its resonant frequency to the offset-charge noise. In other words, the energy difference between the first excited and the ground states has little dependence on the offset charge in the transmon regime $E_{J}/E_{C}\gg 1$. It should be noted that the eigenenergies of higher excited states depend on the offset charge Blais et al. (2021). The presence of the chaotic layer in the spectrum of a driven transmon introduces the offset-charge dependence to the dynamics of low-lying qubit states indeed Cohen et al. (2023). Consequently, the locations of resonances to the higher excited states are expected to depend on the offset-charge noise.
@@
Figure 5: (Color online) Photon-number dependence of the cavity frequency with the offset charges from 00 to $0.5$ for (a) the ground-state ladder |$\ket{\overline{g},\overline{n}}$ and (b) the excited-state ladder |$\ket{\overline{e},\overline{n}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2})$ for the system parameters. For the cases $N_{g}=0.1$ in the ground-state ladder and $N_{g}=0.3$ in the excited-state ladder, the energy window $\delta$ is set to $1.5\times 10^{-2}\hbar\omega_{c}$. For the other cases, we set $\delta$ to $1.0\times 10^{-2}\hbar\omega_{c}$.
@@ VI Summary
In this paper, we proposed a new method to label the eigenstates of a qubit-cavity system based on the continuity of qubit occupancy. Starting from the eigenstate without photon occupation, the system is expected to follow the labeled states under adiabatic injection of photons into the cavity. Besides, the effective resonant frequency of the cavity with photon occupation can be estimated from the eigenenergies of the labeled states. Our proposed method finds candidates for the next labeled state based on their eigenenergies. Then, one selects the next labeled state that makes the change of the qubit occupancy as small as possible. Comparing the obtained photon-number dependence of the cavity frequency and the evolution under cavity driving, we confirmed that our proposed scheme can be applied to a broader situation compared to the previous approach Shillito et al. (2022) in the sense that our proposed scheme can access a case where a system passes through resonant transitions that the previous approach cannot avoid.
With the proposed method, we investigated the offset-charge dependence of the cavity frequency in the transmon-cavity system. In the system we studied, the location of the resonance fluctuates depending on the offset charge. Even though the qubit resonant frequency of the transmon is robust to the charge noises, the resonance to higher excited states is not. Without controlling the offset charge, it is possible to encounter a resonance to higher excited states with a relatively small photon number $n\sim 10$.
The photon-number dependence of the cavity frequency and the location of the resonance to higher excited states are essential information for designing superconducting qubits. Our proposed method is applicable to other superconducting qubits such as the fluxonium Manucharyan et al. (2009) or qubits with $d$–wave superconductors Patel et al. (2024); Brosco et al. (2024). Their robustness to the charge noise, including effects from higher excited states, can be evaluated with a labeling method. A stable labeling method is essential in assessing the robustness of existing or new high-performance superconducting qubits.